find the sum of first n terms of a series whose nth term can be given by
tn=n/(n4+4)

Grade:12

1 Answers

Vikas TU
5 years ago
To find the sum:
First term is given by , put n=1,
 
find t2,put n=2,
d= t2-t1=1/10-1/5=-1/10
sum of first n terms =n/2(2a+(n-1)d)
                   =n/2(2(1/5)+(n-1)(-1/10))
                   =(5n-4n2)/20.
So this would be the sum of the first n terms from the give tn.